OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1). y + 3 = −5(x − 2) y − 2 = −5(x + 3) y + 3 = −1/5(x − 2) y − 2 = −1/5(x + 3)

OpenStudy (goformit100):

I think its y + 3 = −1/5(x − 2)

OpenStudy (anonymous):

slope= (1-2)/ (2-(-3)) = -1/5 y-2 /x-(-3) = -1/5 y-2 =(-1/5)(x+3)